OpenStudy (anonymous):

Does anyone know how to find the tangent and normal line of y=6x^-1 p(3,2)

OpenStudy (anonymous):

the equation of the tangent line to y = f(x) at the point (x1,y1): \[y=m(x-x1)+y1\] where m is the gradient (slope) at (x1,y1) m=\[\frac{ dy }{ dx }=-6x^{-2}\] at x=3 \[m=\frac{ dy }{ dx }=\frac{ -2 }{ 3 }\] so the equation of the tangent line : \[y=\frac{ -2 }{ 3 }(x-3)+2\] \[y =\frac{ -2 }{ 3}x+4\] the equation of the normal line to y = f(x) at (x1,y1) : \[y=\frac{ -1 }{ m}(x-x1)+y1\] \[y=\frac{ -1 }{ \frac{ -2 }{ 3 } }(x-3)+2\] \[y=\frac{ 3 }{ 2 }x-\frac{ -5 }{ 2}\]